OpenStudy (anonymous):

in the equation 2^120 x 10^-21 M^3 = 1.3 x 10^15 M^3 how did the get the 15 exponent!? I am totally lost.

OpenStudy (anonymous):

2^120 = 1.33 * 10^36 when multiplying exponents, you add the power. (substract when dividing) 10^36 + 10^-21 = 10^(36-21) = 10^15

OpenStudy (anonymous):

Euler271 :) THANK YOU!!!!!